OpenStudy (anonymous):

Lily invested $750 in a savings account and earned $225 in interest at the end of 5 years. What was the interest rate? A. 5% B. 5.5% C. 6% D. 6.5%

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

i think it is A: hope it helps

OpenStudy (the_fizicx99):

It being the formula R = I/p*t R = 225/750*5 R = 225/3750 R = 0.06 ---> 6%

OpenStudy (the_fizicx99):

The original formula is I = PRT 225 = 750 * 0.06 * 5 225 = 45 * 5 225 = 225

OpenStudy (anonymous):

np

OpenStudy (anonymous):

so C, again